I1. The pressure of a gas is given by Pwhere P is in atmosphere and V is in litres. If the gas expands from 20 to 60 L and undergocs an increase in internal energy of 225 cal. How much hcat will be absorbed during the process? 12. 5 moles of an ideal gas was initially at 315 K and 20 atm. The expansion of gas takes place adiabatically when the external pressure is reduced to 7 atm. What will be the final temperature and volume? Also calculate the work done during the process, given that Cp -8.58 cal/mol/ C.


can u show the 11th and 12th questions
Answer #1

11) Given V1=20 L V​​2=60 L

P=(15/V) where P is in atm and V is in Liter.

Internal Energy, Δυ--225cal 941.805/

From 1st law of thermodynamics:

Delta U=q+w ......(1)

Where 60 PdT (15/V)dV-16.48/ u1 20

using equation (1) to fi​nd absorbed heat during this process as:

q= (941.850 - 16.48) J= 925.37J

12) No. of moles= 5 , initial temperature and pressure are

T1= 315 K and P1= 20 atm

Using Ideal gas law:

PV=nRT so, V1= (nRT1​/P1) = (5×8.314×315) /(20×101325) = 6.46×10-3 m3= 6.46 L

Since, the system is adiabatic so, it gives relation such as:

P_{1}^{1-gamma }T_{1}^{gamma }=P_{2}^{1-gamma }T_{2}^{gamma }   

T_{2}=T_{1}(P_{1}/P_{2}) ^{(1-gamma )/gamma }.... .(1)

Where P2= 7 atm γ = Cp/C,-Cp/(Cp-R) = 35.914/27.6 1.3

Putting all the values in equation (1) to find final temperature as:

T2 -315(20/7)(1-1.3/1.3 247.2K

Simarly we will find final volume as:

½ = Iİ (Tİ/h)1/(1-1)ー6.46(315/247.2)1/(1.3-1)-14.48し

And work done during this process is:

W=Pext(V2 - V1) = (7×101325) ×(14.48 - 6.46) ×10-3= 5688.4J=5.69KJ